the area of a rectangle is 54 yd^2,and the length of the rectangle is 3 yd more than twice the width. Find the dimensions of the rectangle.
Answer by ikleyn(52756)

.
the area of a rectangle is 54 yd^2,and the length of the rectangle is 3 yd more than twice the width.
Find the dimensions of the rectangle.
~~~~~~~~~~~~~~~

If the width is w yards (now unknown), then the length is (2w+3) yards,

and the area equation is

     w*(2w+3) = 54.      (1)


I will solve it mentally.  For it, I multiply both sides by 2

    2w*(2w+3) = 108.


What are two positive integer numbers that differ by 3 and produce 108, when multiplied ?


You can easy guess them mentally: 9 and 12.


So, 2w = 9;  hence, w = 4.5.


ANSWER.  The dimensions of the rectangle are 4.5 yards (the width) and 12 yards (the length).


CHECK.   4.5*12 = 54 square yards.    ! Correct !

Solved     //     (mentally).

----------------

Alternatively,  you can solve equation  (1)  as a quadratic equation, using routine methods,
such as a quadratic formula or factoring.